# Random generation of universal cycles and de Bruijn sequences

### v1

@@ 4.6. Orientable sequences
An orientable sequence is cyclic sequence such that each length-$n$ substring occurs at most once in either direction. For example, a maximum-length orientable sequence for $n=5$ and $k=2$ is $001101$. A universal cycle for $\mathbf{OS}_{k}(n)$ is known to be an orientable sequence with asymptotically optimal length [6, 3]. A formula for $OS_{k}(n)$, is given in [9, 14]. Generating a random string from $\mathbf{OS}_{k}(n)$ does not appear to be a trivial matter. However, by randomly generating $k$-ary strings with rejection, on average only two random strings need to be generated to obtain a string in $\mathbf{OS}_{k}(n)$ as $n$ gets large. Thus, the expected time to generate a random edge in $G(\mathbf{OS}_{k}(n))$ is $\Theta(n)$.
@@
Orientable sequences with asymptotically optimal length can be constructed in $\mathcal{O}(n)$ time per symbol using $\mathcal{O}(n)$ space [14]; in the binary case, they can be constructed in $\mathcal{O}(1)$-amortized time per bit using $\mathcal{O}(n^{2})$ space.

### v3

@@ 4.6. Orientable sequences
An orientable sequence is cyclic sequence such that each length-$n$ substring occurs at most once in either direction. For example, a maximum-length orientable sequence for $n=5$ and $k=2$ is $001101$. A universal cycle for $\mathbf{OS}_{k}(n)$ is known to be an orientable sequence with asymptotically optimal length [BM, kary-orientable]. A formula for $OS_{k}(n)$, is given in [Dai, korient]. Generating a random string from $\mathbf{OS}_{k}(n)$ does not appear to be a trivial matter. However, by randomly generating $k$-ary strings with rejection, on average only two random strings need to be generated to obtain a string in $\mathbf{OS}_{k}(n)$ as $n$ gets large. Thus, the expected time to generate a random edge in $G(\mathbf{OS}_{k}(n))$ is $\Theta(n)$.
@@
Orientable sequences with asymptotically optimal length can be constructed in $\mathcal{O}(n)$ time per symbol using $\mathcal{O}(n)$ space [korient]; in the binary case, they can be constructed in $\mathcal{O}(1)$ amortized time per bit using $\mathcal{O}(n^{2})$ space.